OpenStudy (anonymous):

is x + 10 a factor of the function f (x) = x^3 - 75x + 250?

OpenStudy (anonymous):

try algebraic long division. If you dont get a remainder then it is factor :)

OpenStudy (anonymous):

hmm can you explain further?

OpenStudy (anonymous):

try that. ANd i will add if you get no remainder left then it is a factor

OpenStudy (anonymous):

i still dont get it at all

OpenStudy (anonymous):

can you do a walkthrough with my example?

jimthompson5910 (jim_thompson5910):

hint: if x-p is a factor of f(x), then f(p) = 0 this is the remainder theorem

OpenStudy (anonymous):

can someone do a step by step walkthrough of my problem please

jimthompson5910 (jim_thompson5910):

x + 10 is the same as x - (-10)

jimthompson5910 (jim_thompson5910):

match x - (-10) up with x - p and you'll see that p = -10

jimthompson5910 (jim_thompson5910):

so you just need to see if f(-10) = 0 is true or not

OpenStudy (anonymous):

not true

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

you need to show why it's not true

jimthompson5910 (jim_thompson5910):

plug x = -10 into f (x) = x^3 - 75x + 250 to see what you get

OpenStudy (anonymous):

-10 cubed is -1000 correct?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

ok then the function would equal

OpenStudy (anonymous):

-1500?

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

you made an error

jimthompson5910 (jim_thompson5910):

try it again

OpenStudy (anonymous):

-1000+750+250= 0 so 0 is the answer

jimthompson5910 (jim_thompson5910):

so because f(-10) = 0, this means that x - (-10) or x + 10 is a factor

OpenStudy (anonymous):

so it is a factor because the remainder is zero?

OpenStudy (anonymous):

or is it a factor because the remainder is NOT zero?

jimthompson5910 (jim_thompson5910):

it's a factor because the remainder is 0

jimthompson5910 (jim_thompson5910):

that's exactly what the factor theorem says above

OpenStudy (anonymous):

thank you so much man!

OpenStudy (anonymous):

hey can you help me with a couple more?

OpenStudy (anonymous):

they deal with remainder questions in long division

jimthompson5910 (jim_thompson5910):

sure, one more